OpenStudy (anonymous):

Write the sentence as an equation. Choose a variable to use. The width of a rectangle is half the length, and the perimeter is 32.

OpenStudy (anonymous):

W = 1/2 L 2W + 2L = 32 Then Solve for one variable

OpenStudy (anonymous):

l=length w=width 2l + 2w = perimeter = 32 2w=l (because the width is half the length) Therefore: 2l+l=32 3l=32 l=32/3= 10.67 (2d.p.) So, l=10.67, and w=10.67/2=5.33

OpenStudy (anonymous):

The width of a rectangle is half the length, and the perimeter is 32.